OpenStudy (anonymous):

Let f(x) = x + 7 and g(x) = x – 2. What is f(x) • g(x)?

OpenStudy (ash2326):

We have \[f(x)=(x+7)\] \[g(x)=x-2\] \[f(x).g(x)=(x+7)\times(x-2)\] Can you take it from here?

OpenStudy (anonymous):

Yes thank you. I've got the answer.